# OmniCT: Towards a Unified Slice-Volume LVLM for Comprehensive CT Analysis

### v1

@@ Abstract
Computed Tomography (CT) is one of the most widely used and diagnostically information-dense imaging modalities, covering critical organs such as the heart, lungs, liver, and colon. Clinical interpretation relies on both slice-driven local features (e.g., sub-centimeter nodules, lesion boundaries) and volume-driven spatial representations (e.g., tumor infiltration, inter-organ anatomical relations). However, existing Large Vision–Language Models (LVLMs) remain fragmented in CT slice versus volumetric understanding: slice-driven LVLMs show strong generalization but lack cross-slice spatial consistency, while volume-driven LVLMs explicitly capture volumetric semantics but suffer from coarse granularity and poor compatibility with slice inputs. The absence of a unified modeling paradigm constitutes a major bottleneck for the clinical translation of medical LVLMs. We present OmniCT, a powerful unified slice–volume LVLM for CT scenarios, which makes three contributions: (i) Spatial Consistency Enhancement (SCE): volumetric slice composition combined with tri-axial positional embedding that introduces volumetric consistency, and an MoE hybrid projection enables efficient slice–volume adaptation; (ii) Organ-level Semantic Enhancement (OSE): segmentation and ROI localization explicitly align anatomical regions, emphasizing lesion- and organ-level semantics; (iii) MedEval-CT: the largest slice–volume CT dataset and hybrid benchmark integrates comprehensive metrics for unified evaluation. OmniCT consistently outperforms existing methods with a substantial margin across diverse clinical tasks and satisfies both micro-level detail sensitivity and macro-level spatial reasoning. More importantly, it establishes a new paradigm for cross-modal medical imaging understanding.
@@ 1 Introduction
Large Vision–Language Models (LVLMs) have become a cornerstone of multi-modal artificial intelligence, demonstrating strong cross-modal representation and reasoning capabilities in both image understanding (Qiu et al., 2023; Li et al., 2024a; Zhu et al., 2025; Bai et al., 2025b) and video perception (Lin et al., 2023a; Li et al., 2024a; Zhang et al., 2025a; Yuan et al., 2025). Benefiting from large-scale pre-training and modality alignment, LVLMs achieve remarkable performance in open-domain tasks (Qiu et al., 2022; Yue et al., 2024; Fu et al., 2025), excelling in both generation and reasoning. These advances establish LVLMs as a universal paradigm for unified vision–language modeling, where the joint modeling of 2D and 3D modalities has emerged as a key design principle.
11footnotetext: Equal contributions. The work was done during Tianwei’s internship at DAMO Academy.22footnotetext: Corresponding author.
@@
In recent years, the potential of LVLMs in medical imaging has received increasing attention, with exploration in radiological imaging being particularly notable (Wu et al., 2025; Xu et al., 2025). However, most existing methods are tailored to process either CT slices (Chen et al., 2024a; Lin et al., 2025) or volumetric data (Bai et al., 2024; Hamamci et al., 2024c), with limited focus on cooperative processing. Slice-driven models leverage large-scale 2D pre-training to achieve strong vision–language alignment and perform well in tasks such as lesion detection and radiology report description, yet they fail to capture cross-slice spatial consistency. In contrast, volume-driven models explicitly model voxel-level spatial structures, offering advantages in holistic spatial representation and organ-level reasoning. Nevertheless, these models often lack sensitivity to fine-grained abnormalities and boundary morphology, and their architectures are difficult to adapt to slice-level tasks, limiting their applicability across diverse medical scenarios. This persistent dichotomy between slice and volume modeling constitutes a major bottleneck in the development of medical LVLMs.
@@
Among various medical imaging modalities, CT is one of the most widely used and dense in information, with hundreds of million performed each year globally. CT can cover critical organs such as the heart, lungs, liver, and colon, and is widely applied in essential tasks including disease screening (Hu et al., 2025), lesion assessment (Li et al., 2025b; Shui et al., 2025), and tumor staging (Bassi et al., 2025; Qiu et al., 2025). Its diagnostic process relies on both slice-level local imaging cues, such as sub-centimeter pulmonary nodules or hepatic lesion boundaries, and volume-level spatial–topological representations, such as tumor infiltration ranges or inter-organ anatomical relationships. Modeling along a single dimension alone cannot meet these dual requirements. Therefore, integrating the complementary strengths of 2D and 3D modeling within a unified framework is not only a central scientific challenge for CT understanding but also an inevitable step toward the clinical translation of medical LVLMs.
@@
We propose OmniCT (see Fig. 1), a powerful unified slice–volume LVLM for CT-centric understanding, which preserves the cross-modal alignment and generalization strengths of 2D models while integrating the spatial structural awareness of 3D models. To bridge the modality gap between slice and volume representations, we introduce a Spatial Consistency Enhancement (SCE) strategy. Unlike generic LVLMs that rely on frame sampling or key-frame stacking strategies (Xu et al., 2024; Li et al., 2024b; Huang et al., 2024), SCE performs Volumetric Slice Composition (VSC) by structurally combining adjacent slices along the channel dimension into locally consistent volumetric units, thereby retaining contextual spatial transitions. It further incorporates a Tri-axial Positional Encoding (TPE), which injects 3D positional encodings into visual representations to enable volumetric awareness while maintaining compatibility with slice-based inputs. In addition, a MoE Hybrid Projection (MHP) dynamically aligns slice and volume features within a shared representation space, ensuring semantic unification with the Large Language Models (LLMs). Overall, SCE injects robust 2D/3D spatial priors while achieving a balance between efficiency and adaptability.
@@
In clinical diagnosis, image interpretation is performed at the organ level, where observations and lesion localization are conducted within this scope (Shui et al., 2025). Building on this clinical requirement, we propose Organ-level Semantic Enhancement (OSE). OSE performs task-guided anatomical region localization, explicitly projecting critical organ regions into the token representation space and fusing them with global visual context, thereby embedding organ-centric semantics into the representation. It then applies an adaptive aggregation to compress long-sequence representations: this mechanism preserves overall information coverage while adaptively magnifying smaller organ regions and compressing larger ones, thus highlighting the most diagnostically relevant structures. In this way, OSE explicitly incorporates region priors with high task-relevant semantic load while improving the relevance and interpretability of models in clinical tasks.
@@
Existing medical benchmarks (Hu et al., 2024; Yue et al., 2024; Yamagishi et al., 2025) often adopt multi-modality designs to evaluate the general capability of LVLMs, yet they fall short in task alignment and clinical representativeness for CT interpretation. To address this gap, we introduce MedEval-CT, the first holistic evaluation framework dedicated to CT images. At the data level, MedEval-CT-Dataset consolidates 1.7M slice-driven and volume-driven VQA samples across 7 clinical task types, establishing the largest CT resource to date (Fig. 1 (a)). At the benchmark level, MedEval-CT-Bench organizes hybrid evaluations along clinical problem types and organ distributions. At the toolkit level, MedEval-CT-Factory standardizes input handling, feature construction, and multi-dimensional metrics, supporting statistical, semantic, and LLM-based evaluations. Collectively, MedEval-CT institutionalizes fairness and comparability in medical LVLM evaluation, while providing a scalable foundation for larger and more complex clinical scenarios.
@@ 2.2 Organ-level Semantic Enhancement
Anatomical region localization. Given the visual token representation $\hat{\mathcal{F}}\in\mathbb{R}^{L\times d_{h}}$ produced by SCE, we perform region-wise selection based on spatial priors of the target organ $o$. The organ mask is denoted as $\mathcal{M}_{o}\in\mathbb{R}^{D\times H\times W}$, including 117 anatomical structures, which is generated by TotalSegmentor (Wasserthal et al., 2023). This mask of $D\times H\times W$ is mapped to the token size by leveraging the scaling relationship between pixels and vision tokens, resulting in the organ-specific subset: $\hat{\mathcal{F}}_{o}=\hat{\mathcal{F}}[\hat{\mathcal{M}}_{o}],$ where $\hat{\mathcal{M}}_{o}]$ denotes mask-based indexing for token selection. $\hat{\mathcal{F}}_{o}$ represents the selected tokens of size $L_{o}\times d_{h}$ for organ $o$ by the organ mask $\hat{\mathcal{M}}_{o}$.
@@
where $L_{c}$ denotes the fixed number of aggregated tokens compressed from $L_{o}$. This aggregation not only reduces token redundancy but also introduces a “magnification effect” for small organs, enhancing fine-grained lesion features. Simultaneously, it applies a “compression effect” to large organs or global regions, effectively minimizing redundancy and preserving essential information.
@@ 3.1 MedEval-CT
MedEval-CT-Dataset. We introduce MedEval-CT-Dataset, the largest unified CT imaging resource to date, comprising over 1.7 million VQA samples from 170,280 independent 3D volumes and 327,063 standalone 2D slices. The two subsets are collected from distinct, non-overlapping sources, ensuring that the 2D slices are neither extracted nor sampled from the 3D volumes. To fully leverage the rich spatial structure and clinical semantics of volumetric CT data, each 3D volume is annotated with multiple VQA instances covering diverse diagnostic perspectives, whereas the 2D subset mainly supports slice-level interpretation with typically one question per slice. Overall, the dataset enables comprehensive evaluation of both 2D clinical understanding and 3D volumetric perception, accounting for 16.3% and 83.7% of the data, respectively. As shown in Fig. 3, the dataset is systematically partitioned across three dimensions: task types, clinical categories, and organs, enabling multi-faceted evaluation of LVLMs. For task types, it spans seven medical VQA scenarios, from structured to open-ended tasks: (1) Free-form QA (38.0%) (2) Multiple Choice (18.0%) (3) Short Answers (13.7%) (4) Report Generation (11.5%) (5) Caption (7.2%) (6) Dialogue (6.2%) and (7) Judgment (5.5%). Clinical categories reflect increasing difficulty, progressing from basic anatomical recognition to expert-level reasoning: (1) General Imaging Recognition (GIR, 21.9%) (2) Medical Abnormality Identification (MAI, 23.6%) (3) Advanced Imaging Interpretation (AII, 34.5%) and (4) Clinical Reasoning and Decision (CRD, 19.9%). Organ-wise, it covers lungs (32.4%), vessels (8.2%), heart (6.6%), liver (4.9%), kidneys (4.3%), and additional regions like spine, trachea, and esophagus, ensuring robust anatomical diversity. Overall, MedEval-CT surpasses existing datasets in scale and granularity, providing high-resolution distributions across tasks, clinical expertise, and organs to advance the development of LVLMs for CT imaging. Data sources and other details are presented in Table C and Appendix E.
@@
MedEval-CT-Factory. We introduce MedEval-CT-Factory, an institutionalized evaluation factory designed to address the heterogeneity of inputs, features, and outputs in medical LVLMs. At the input level, the Factory standardizes diverse CT data formats, including DICOM, NIfTI, arrays, and slice sequences, enabling seamless 2D/3D processing. At the feature level, it unifies model inputs (single images, multi-slice sequences, videos, or volumes) via frame sampling, resampling, and projection strategies. At the output level, it provides a multi-layer evaluation protocol, ranging from statistical metrics (BLEU (Papineni et al., 2002), ROUGE (Lin, 2004)), to semantic metrics (BERTScore (Zhang et al., 2019), embedding similarity (Zhang et al., 2025b)), and further to LLM-based evaluation simulating clinical reasoning. Overall, MedEval-CT-Factory streamlines complex engineering workflows into a standardized framework, ensuring comparability across models. Serving as the fourth pillar of the MedEval-CT paradigm alongside the Dataset, Bench, and Data Engine, it is designed as a toolbox to enhance both the efficiency and fairness of LVLM evaluation in the CT domain. (The framework and details of MedEval-CT-Factory are shown in Appendix G).
@@ 4 Experiments
Baseline Comparisons. To comprehensively evaluate the performance of OmniCT against existing open-source medical LVLMs as well as strong general-purpose LVLMs, we select a diverse set of baseline models that systematically cover both the general-to-medical spectrum and the 2D-to-3D setting. For the 2D slice-based benchmarks, we include general-purpose LVLMs such as InternVL3 (Zhu et al., 2025), Qwen2.5-VL (Bai et al., 2025b), and GPT-5 (Wang et al., 2025), together with representative medical-domain models including HealthGPT (Lin et al., 2025), HuatuoGPT-Vision-Qwen2.5 (Chen et al., 2024b), MedGemma-4B-IT (Sellergren et al., 2025), MedVLM-R1-2B (Pan et al., 2025), Lingshu (Xu et al., 2025), and RadFM (Wu et al., 2025). For the 3D volume-based benchmarks, we compare OmniCT with strong general LVLMs (MiniCPM-V-4.5 (Yu et al., 2025), Qwen2.5-VL, and GPT-5) as well as specialized 3D medical LVLMs, including M3D-LaMed (Bai et al., 2024), CT-CHAT (Hamamci et al., 2024a), and RadFM. The Data Details, Model Details, and Implementation Details are shown in Appendix E.

### v2

# arXiv:2602.16110v2[cs.CV] 01 Mar 2026

@@ Abstract
Computed Tomography (CT) is one of the most widely used and diagnostically information-dense imaging modalities, covering critical organs such as the heart, lungs, liver, and colon. Clinical interpretation relies on both slice-driven local features (e.g., sub-centimeter nodules, lesion boundaries) and volume-driven spatial representations (e.g., tumor infiltration, inter-organ anatomical relations). However, existing Large Vision–Language Models (LVLMs) remain fragmented in CT slice versus volumetric understanding: slice-driven LVLMs show strong generalization but lack cross-slice spatial consistency, while volume-driven LVLMs explicitly capture volumetric semantics but suffer from coarse granularity and poor compatibility with slice inputs. The absence of a unified modeling paradigm constitutes a major bottleneck for the clinical translation of medical LVLMs. We present OmniCT, a powerful unified slice–volume LVLM for CT scenarios, which makes three contributions: (i) Spatial Consistency Enhancement (SCE): volumetric slice composition combined with tri-axial positional embedding that introduces volumetric consistency, and an MoE hybrid projection enables efficient slice–volume adaptation; (ii) Organ-level Semantic Enhancement (OSE): segmentation and ROI localization explicitly align anatomical regions, emphasizing lesion- and organ-level semantics; (iii) MedEval-CT: the largest slice–volume CT dataset and hybrid benchmark integrates comprehensive metrics for unified evaluation. OmniCT consistently outperforms existing methods with a substantial margin across diverse clinical tasks and satisfies both micro-level detail sensitivity and macro-level spatial reasoning. More importantly, it establishes a new paradigm for cross-modal medical imaging understanding. Our project is available at https://github.com/ZJU4HealthCare/OmniCT.
@@ 1 Introduction
Large Vision–Language Models (LVLMs) have become a cornerstone of multi-modal artificial intelligence, demonstrating strong cross-modal representation and reasoning capabilities in both image understanding (qiu2023psvt; li2024llava; zhu2025internvl3; bai2025qwen2) and video perception (lin2023video; li2024llava; zhang2025videollama; yuan2025videorefer). Benefiting from large-scale pre-training and modality alignment, LVLMs achieve remarkable performance in open-domain tasks (qiu2022learning; yue2024mmmu; fu2025video), excelling in both generation and reasoning. These advances establish LVLMs as a universal paradigm for unified vision–language modeling, where the joint modeling of 2D and 3D modalities has emerged as a key design principle.
@@
In recent years, the potential of LVLMs in medical imaging has received increasing attention, with exploration in radiological imaging being particularly notable (wu2025towards; xu2025lingshu). However, most existing methods are tailored to process either CT slices (chen2024huatuogpt; lin2025healthgpt) or volumetric data (bai2024m3d; hamamci2024developing), with limited focus on cooperative processing. Slice-driven models leverage large-scale 2D pre-training to achieve strong vision–language alignment and perform well in tasks such as lesion detection and radiology report description, yet they fail to capture cross-slice spatial consistency. In contrast, volume-driven models explicitly model voxel-level spatial structures, offering advantages in holistic spatial representation and organ-level reasoning. Nevertheless, these models often lack sensitivity to fine-grained abnormalities and boundary morphology, and their architectures are difficult to adapt to slice-level tasks, thereby limiting their applicability across diverse medical scenarios. This persistent dichotomy between slice and volume modeling constitutes a major bottleneck in the development of medical LVLMs.
@@
Among various medical imaging modalities, CT is one of the most widely used and dense in information, with hundreds of million performed each year globally. CT can cover critical organs such as the heart, lungs, liver, and colon, and is widely applied in essential tasks including disease screening (hu2025ai), lesion assessment (LiYin_More_MICCAI2025; shui2025large), and tumor staging (bassi2025radgpt; qiu2025bridging). Its diagnostic process relies on both slice-level local imaging cues, such as sub-centimeter pulmonary nodules or hepatic lesion boundaries, and volume-level spatial–topological representations, such as tumor infiltration ranges or inter-organ anatomical relationships. Modeling along a single dimension alone cannot meet these dual requirements. Therefore, integrating the complementary strengths of 2D and 3D modeling within a unified framework is not only a central scientific challenge for CT understanding but also an inevitable step toward the clinical translation of medical LVLMs.
@@
We propose OmniCT (see Fig. 1), a powerful unified slice–volume LVLM for CT-centric understanding, which preserves the cross-modal alignment and generalization strengths of 2D models while integrating the spatial structural awareness of 3D models. To bridge the modality gap between slice and volume representations, we introduce a Spatial Consistency Enhancement (SCE) strategy. Unlike generic LVLMs that rely on frame sampling or key-frame stacking strategies (xu2024pllava; li2024llama; huang2024lita), SCE performs Volumetric Slice Composition (VSC) by structurally combining adjacent slices along the channel dimension into locally consistent volumetric units, thereby retaining contextual spatial transitions. It further incorporates a Tri-axial Positional Encoding (TPE), which injects 3D positional encodings into visual representations to enable volumetric awareness while maintaining compatibility with slice-based inputs. In addition, a MoE Hybrid Projection (MHP) dynamically aligns slice and volume features within a shared representation space, ensuring semantic unification with the Large Language Models (LLMs). Overall, SCE injects robust 2D/3D spatial priors while achieving a balance between efficiency and adaptability.
@@
In clinical diagnosis, image interpretation is performed at the organ level, where observations and lesion localization are conducted within this scope (shui2025large). Building on this clinical requirement, we propose Organ-level Semantic Enhancement (OSE). OSE performs task-guided anatomical region localization, explicitly projecting critical organ regions into the token representation space and fusing them with global visual context, thereby embedding organ-centric semantics into the representation. It then applies an adaptive aggregation to compress long-sequence representations: this mechanism preserves overall information coverage while adaptively magnifying smaller organ regions and compressing larger ones, thus highlighting the most diagnostically relevant structures. In this way, OSE explicitly incorporates region priors with high task-relevant semantic load while improving the relevance and interpretability of models in clinical tasks.
@@
Existing medical benchmarks (hu2024omnimedvqa; yue2024mmmu; yamagishi2025radfig) often adopt multi-modality designs to evaluate the general capability of LVLMs, yet they fall short in task alignment and clinical representativeness for CT interpretation. To address this gap, we introduce MedEval-CT, the first holistic evaluation framework dedicated to CT images. At the data level, MedEval-CT-Dataset consolidates 1.7M slice-driven and volume-driven VQA samples across 7 clinical task types, establishing the largest CT resource to date (Fig. 1 (a)). At the benchmark level, MedEval-CT-Bench organizes hybrid evaluations along clinical problem types and organ distributions. At the toolkit level, MedEval-CT-Factory standardizes input handling, feature construction, and multi-dimensional metrics, supporting statistical, semantic, and LLM-based evaluations. Collectively, MedEval-CT institutionalizes fairness and comparability in medical LVLM evaluation, while providing a scalable foundation for larger and more complex clinical scenarios.
@@ 2.2 Organ-level Semantic Enhancement
Anatomical region localization. Given the visual token representation $\hat{\mathcal{F}}\in\mathbb{R}^{L\times d_{h}}$ produced by SCE, we perform region-wise selection based on spatial priors of the target organ $o$. The organ mask is denoted as $\mathcal{M}_{o}\in\mathbb{R}^{D\times H\times W}$, including 117 anatomical structures, which is generated by TotalSegmentor (wasserthal2023totalsegmentator). This mask of $D\times H\times W$ is mapped to the token size by leveraging the scaling relationship between pixels and vision tokens, resulting in the organ-specific subset: $\hat{\mathcal{F}}_{o}=\hat{\mathcal{F}}[\hat{\mathcal{M}}_{o}],$ where $\hat{\mathcal{M}}_{o}]$ denotes mask-based indexing for token selection. $\hat{\mathcal{F}}_{o}$ represents the selected tokens of size $L_{o}\times d_{h}$ for organ $o$ by the organ mask $\hat{\mathcal{M}}_{o}$.
@@
where $L_{c}$ denotes the fixed number of aggregated tokens compressed from $L_{o}$. This aggregation not only reduces token redundancy but also introduces a “magnification effect" for small organs, enhancing fine-grained lesion features. Simultaneously, it applies a “compression effect" to large organs or global regions, effectively minimizing redundancy and preserving essential information.
@@ 3.1 MedEval-CT
MedEval-CT-Dataset. We introduce MedEval-CT-Dataset, the largest unified CT imaging resource to date, comprising over 1.7 million VQA samples from 170,280 independent 3D volumes and 327,063 standalone 2D slices. The two subsets are collected from distinct, non-overlapping sources, ensuring that the 2D slices are neither extracted nor sampled from the 3D volumes. To fully leverage the rich spatial structure and clinical semantics of volumetric CT data, each 3D volume is annotated with multiple VQA instances covering diverse diagnostic perspectives, whereas the 2D subset mainly supports slice-level interpretation with typically one question per slice. Overall, the dataset enables comprehensive evaluation of both 2D clinical understanding and 3D volumetric perception, accounting for 16.3% and 83.7% of the data, respectively. As shown in Fig. 3, the dataset is systematically partitioned across three dimensions: task types, clinical categories, and organs, enabling multi-faceted evaluation of LVLMs. For task types, it spans seven medical VQA scenarios, from structured to open-ended tasks: (1) Free-form QA (38.0%) (2) Multiple Choice (18.0%) (3) Short Answers (13.7%) (4) Report Generation (11.5%) (5) Caption (7.2%) (6) Dialogue (6.2%) and (7) Judgment (5.5%). Clinical categories reflect increasing difficulty, progressing from basic anatomical recognition to expert-level reasoning: (1) General Imaging Recognition (GIR, 21.9%) (2) Medical Abnormality Identification (MAI, 23.6%) (3) Advanced Imaging Interpretation (AII, 34.5%) and (4) Clinical Reasoning and Decision (CRD, 19.9%). Organ-wise, it covers lungs (32.4%), vessels (8.2%), heart (6.6%), liver (4.9%), kidneys (4.3%), and additional regions like spine, trachea, and esophagus, ensuring robust anatomical diversity. Overall, MedEval-CT surpasses existing datasets in scale and granularity, providing high-resolution distributions across tasks, clinical expertise, and organs to advance the development of LVLMs for CT imaging. Data sources and other details are presented in Table 8 and Appendix E.
@@
MedEval-CT-Factory. We introduce MedEval-CT-Factory, an institutionalized evaluation factory designed to address the heterogeneity of inputs, features, and outputs in medical LVLMs. At the input level, the Factory standardizes diverse CT data formats, including DICOM, NIfTI, arrays, and slice sequences, enabling seamless 2D/3D processing. At the feature level, it unifies model inputs (single images, multi-slice sequences, videos, or volumes) via frame sampling, resampling, and projection strategies. At the output level, it provides a multi-layer evaluation protocol, ranging from statistical metrics (BLEU (papineni2002bleu), ROUGE (lin2004rouge)), to semantic metrics (BERTScore (zhang2019bertscore), embedding similarity (qwen3embedding)), and further to LLM-based evaluation simulating clinical reasoning. Overall, MedEval-CT-Factory streamlines complex engineering workflows into a standardized framework, ensuring comparability across models. Serving as the fourth pillar of the MedEval-CT paradigm alongside the Dataset, Bench, and Data Engine, it is designed as a toolbox to enhance both the efficiency and fairness of LVLM evaluation in the CT domain. (The framework and details of MedEval-CT-Factory are shown in Appendix G).
@@ 4 Experiments
Baseline Comparisons. To comprehensively evaluate the performance of OmniCT against existing open-source medical LVLMs as well as strong general-purpose LVLMs, we select a diverse set of baseline models that systematically cover both the general-to-medical spectrum and the 2D-to-3D setting. For the 2D slice-based benchmarks, we include general-purpose LVLMs such as InternVL3 (zhu2025internvl3), Qwen2.5-VL (bai2025qwen2), and GPT-5 (wang2025capabilities), together with representative medical-domain models including HealthGPT (lin2025healthgpt), HuatuoGPT-Vision-Qwen2.5 (chen2024huatuogptvisioninjectingmedicalvisual), MedGemma-4B-IT (sellergren2025medgemma), MedVLM-R1-2B (pan2025medvlm), Lingshu (xu2025lingshu), and RadFM (wu2025towards). For the 3D volume-based benchmarks, we compare OmniCT with strong general LVLMs (MiniCPM-V-4.5 (yu2025minicpmv45cookingefficient), Qwen2.5-VL, and GPT-5) as well as specialized 3D medical LVLMs, including M3D-LaMed (bai2024m3d), CT-CHAT (hamamci2024foundation), and RadFM. The Data Details, Model Details, and Implementation Details are shown in Appendix E.
4.1 Main Experiments